### v1

@@ Abstract
This paper presents the first application of Gaussian Mixture Copula Models to the statistical modeling of driving scenarios for the safety validation of automated driving systems. Knowledge of the joint probability distribution of scenario parameters is essential for scenario-based safety assessment, where risk quantification depends on the likelihood of concrete parameter combinations. Gaussian Mixture Copula Models bring together the multimodal expressivity of Gaussian Mixture Models and the flexibility of copulas, enabling separate modeling of marginal distributions and dependencies. We benchmark Gaussian Mixture Copula Models against previously proposed approaches—Gaussian Mixture Models and Gaussian Copula Models—using real-world driving data drawn from scenarios defined in United Nations Regulation No. 157. Our evaluation across 18 million scenario instances demonstrates that Gaussian Mixture Copula Models provide a better fit to the data in terms of both likelihood and Sinkhorn distance. These results suggest that Gaussian Mixture Copula Models are a compelling foundation for future scenario-based validation frameworks.
@@ 1 INTRODUCTION
Achieving safety levels comparable to human drivers requires extensive testing and validation [4]. However, given the impracticality of real-world testing—which would require billions of kilometers of driving data and substantial financial and time investments [5]—simulation has become an indispensable tool for scenario-based assessments. Scenario-based approaches decompose the overall validation task into logical traffic scenarios, such as cut-ins or lead vehicle decelerations [6, 7, 8]. These scenarios are defined by key parameters, such as braking deceleration and inter-vehicular distance, as illustrated in the Deceleration scenario in Fig. 1. By systematically varying these parameters within simulation environments, the space of potential scenarios is explored to identify critical combinations that pose safety risks.
Central to this approach is the ability to determine the joint probability of specific parameter combinations, as this directly underpins the calculation of risk and ensures that simulated scenarios are relevant for real-world operation. ISO 21448 defines risk as “the combination of the probability of occurrence of harm and the severity of that harm” [1, Clause 3.23]. Without precise knowledge of these probabilities, the residual risk cannot be effectively quantified, nor can the most critical scenarios for validation be prioritized [9].
@@
  • Gaussian Mixture Models (GMMs) approximate the joint density with a weighted sum of elliptical Gaussians [12] and have been successfully applied to driving scenarios, such as the highway cut-in risk analysis by Akagi et al. [13]. Nevertheless GMMs are limited in their ability to capture intricate dependencies between parameters due to their reliance on Gaussian assumptions. These assumptions often oversimplify complex relationships and higher-order interactions, restricting their ability to model non-linear dependencies effectively [14].
@@
This work presents the first application of GMCMs to the estimation of joint probability distributions of scenario parameters in driving scenarios. We demonstrate how GMCMs can be used to model complex, multimodal dependencies in safety-critical scenarios, as required by ISO 21448 and UN Regulation No. 157. Our evaluation is based on 18 million real-world driving scenarios and includes a quantitative comparison with GMM and Gaussian Copula Models (GCM) using both log-likelihood and Sinkhorn distance metrics.
@@ A GMM describes the joint density of random variables as a weighted sum of $K$ normal distributions,
where $\alpha^{k}\geq 0$ denotes the weight of the $k$-th component, such that $\sum_{k=1}^{K}\alpha^{k}=1$. $\phi(x_{1},x_{2},\dots x_{d};\bm{\mu}^{k},\Sigma^{k})$ represents a multivariate normal distribution with mean vector $\bm{\mu}^{k}$ and covariance matrix $\Sigma^{k}$. The parameters $\Theta=\{\alpha^{k},\bm{\mu}^{k},\Sigma^{k}\}_{k=1}^{K}$ of a GMM can be efficiently learned by using the expectation-maximization (EM) algorithm [22]. The two-step approach of the EM algorithm guarantees convergence to a local maximum of the likelihood function. As stated before, a major limitation of GMMs is the assumption that each of the $K$ components follows a multivariate normal distribution. This assumption becomes restrictive where the data exhibits non-Gaussian characteristics, such as skewness, heavy-tailed distributions, or intricate dependencies between variables, like this is the case for our scenario parameters.
@@ 3 ESTIMATION OF GMCM PARAMETERS
x $\in\mathbb{V}^{d}=\bigl{[}F^{-1}_{1}(u_{1}),\,F^{-1}_{2}(u_{2}),\,\dots\,F^{-1}_{d}(u_{d})\bigr{]}.$ (8b)
@@
[…]
@@ 4 EXPERIMENT
In this section, we present our experimental study of GMCMs for joint probability estimation of scenario parameters. We demonstrate that GMCMs surpass both GMMs and GCMs in accurately capturing the complex parameter dependencies. Our experiments conducted on real‐world datasets for certification relevant scenarios, confirm the previously reported advantages of GMCMs [17, 21]. They also extend these benefits to the domain of scenario parameter modeling.
@@ 4.1 REAL-WORLD SCENARIO DATASETS
    Deceleration: the ‘other vehicle’ suddenly decelerates in front of the ‘ego vehicle’
@@
For our investigation, we analyzed 14 million and 4 million real-world instances of the Follow Lead Vehicle Emergency Brake and Deceleration scenario, respectively. The data is derived from the Mercedes-Benz customer fleet, which collects information whenever its assistance systems detect a critical traffic situation, provided that customers have given their consent. In such cases, a full report describing the situation is sent to the backend for further analysis. For more details about the data used in this study, refer to [35].
@@ 4.2 RESULTS
In the second step, the MAP procedure for the estimation of the dependencies $\Theta$ among the scenario parameters has been carried out using the Adam Optimizer with the default learning rate of $1E^{-3}$ [37]. The best-fit GMCM and GMM have been obtained on both datasets, wherein the number of components has been chosen to be four. An automated selection using criteria like the Bayesian Information Criterion (BIC) is proposed for future work (see section 5).
To illustrate the achieved expressive power of the GMCM, we visualize the bivariate marginal distribution of LeadVehicle Init HeadwayTime (s) and Ego InitSpeed Ve0 (kph) via density heatmaps. Fig. 4 compares the original data distribution—which appears multimodal with non-Gaussian modes—with the estimated densities from the GMCM, the GMM, and the GCM. Visual inspection of the bivariate density heatmaps suggests that the GMCM most accurately captures the real data distribution.
@@
Unlike the heatmaps, which only allow model evaluation in two dimensions, this observation is quantitatively supported across the full-dimensional space by the Sinkhorn distance and the log-likelihood. Using these metrics, the GMCM also demonstrates improved performance in an analogous evaluation on the Deceleration scenario data set, as shown in Table 3.
Table 3: Log-likelihood and Sinkhorn distance achieved by GMCM, GMM and GCM on our Follow Lead Vehicle Emergency Brake and Deceleration Dataset. The best-performing model for each metric and dataset is shown in bold, illustrating that GMCM consistently achieves a closer fit to the data than GMM or GCM.
Model Log Likelihood Sinkhorn Distance
GMM -7.0914 0.633
GCM -7.1706 0.510
GMCM -6.7923 0.436
@@
Model Log Likelihood Sinkhorn Distance
GMM -9.7469 2.08
GCM -12.0305 11.8
GMCM -9.6185 1.69
@@
The Python implementation we used to achieve the results presented in this section is provided with the paper.
@@ 5 CONCLUSION
GMCMs demonstrated superior performance in all conducted experiments, even without fine-tuning of hyperparameters. This suggests that GMCMs offer a promising alternative to established approaches for probabilistic modeling of scenario parameters. Future work should explore the extent to which performance can be improved by tuning estimation parameters, such as selecting the optimal number of mixture components via the BIC and adjusting convergence criteria [41]. Moreover, GMCMs should be applied to other scenario types defined in UN R157, such as Cut-in and Cut-out. Their performance in higher-dimensional settings should also be evaluated, for example by incorporating parameters related to roadway and environmental conditions. Finally, we restricted our analysis to continuous parameters. How discrete scenario parameters could be integrated into the GMCM framework remains an open question and is left for future work.

### v2

@@ Abstract
This paper presents the first application of Gaussian Mixture Copula Models to the statistical modeling of driving scenarios for the safety validation of automated driving systems. Knowledge of the joint probability distribution of scenario parameters is essential for scenario-based safety assessment, where risk quantification depends on the likelihood of concrete parameter combinations. Gaussian Mixture Copula Models bring together the multimodal expressivity of Gaussian Mixture Models and the flexibility of copulas, enabling separate modeling of marginal distributions and dependencies. We benchmark Gaussian Mixture Copula Models against previously proposed approaches—Gaussian Mixture Models and Gaussian Copula Models—using real-world driving data drawn from scenarios defined in United Nations Regulation No. 157. Our evaluation across approximately 18 million scenario instances demonstrates that Gaussian Mixture Copula Models consistently surpass Gaussian Copula Models and perform better than, or at least comparably to, Gaussian Mixture Models, as measured by both log-likelihood and Sinkhorn distance. These results are promising for the adoption of Gaussian Mixture Copula Models as a statistical foundation for future scenario-based validation frameworks.
@@ 1 INTRODUCTION
Achieving safety levels comparable to human drivers requires extensive testing and validation [4]. However, given the impracticality of real-world testing—which would require billions of kilometers of driving data and substantial financial and time investments [5]—simulation has become an indispensable tool for scenario-based assessments (see also [6, 7]). Scenario-based approaches decompose the overall validation task into logical traffic scenarios, such as cut-ins or lead vehicle decelerations [8, 9, 10]. These scenarios are defined by key parameters, such as braking deceleration and inter-vehicular distance, as illustrated in the Deceleration scenario in Fig. 1. By systematically varying these parameters within simulation environments, the space of potential scenarios is explored to identify critical combinations that pose safety risks.
Central to this approach is the ability to determine the joint probability of specific parameter combinations, as this directly underpins the calculation of risk and ensures that simulated scenarios are relevant for real-world operation. ISO 21448 defines risk as the “combination of the probability of occurrence of harm and the severity of that harm” [1, subclause 3.23]. Without precise knowledge of these probabilities, the residual risk cannot be effectively quantified, nor can the most critical scenarios for validation be prioritized [11].
@@
  • Gaussian Mixture Models (GMMs) approximate the joint density with a weighted sum of elliptical Gaussians [14] and have been successfully applied to driving scenarios, such as the highway cut-in risk analysis by Akagi et al. [15]. Nevertheless, for finite mixtures, GMMs can be limited in practice at capturing intricate dependencies between parameters due to their reliance on Gaussian assumptions. These assumptions often oversimplify complex relationships and higher-order interactions, restricting their ability to model non-linear dependencies effectively [16].
@@
This work presents the first application of GMCMs to the estimation of joint probability distributions of scenario parameters in driving scenarios. We demonstrate how GMCMs can be used to model complex, multimodal dependencies in safety-critical scenarios, as required by ISO 21448 and UN Regulation No. 157. Our evaluation is based on approximately 18 million real-world driving scenarios and includes a quantitative comparison with GMM and Gaussian Copula Models (GCM) using both log-likelihood and Sinkhorn distance metrics.
@@ A GMM describes the joint density of random variables as a weighted sum of $K$ normal distributions,
where $\alpha^{k}\geq 0$ denotes the weight of the $k$-th component, such that $\sum_{k=1}^{K}\alpha^{k}=1$. $\phi(x_{1},x_{2},\dots x_{d};\bm{\mu}^{k},\Sigma^{k})$ represents a multivariate normal distribution with mean vector $\bm{\mu}^{k}$ and covariance matrix $\Sigma^{k}$. The parameters $\Theta=\{\alpha^{k},\bm{\mu}^{k},\Sigma^{k}\}_{k=1}^{K}$ of a GMM can be efficiently learned by using the expectation-maximization (EM) algorithm [24]. The two-step approach of the EM algorithm guarantees convergence to a local maximum of the likelihood function. As stated before, a major limitation of GMMs is the assumption that each of the $K$ components follows a multivariate normal distribution. This assumption can become restrictive in practice when the data exhibit non-Gaussian characteristics, such as skewness, heavy-tailed distributions, or intricate dependencies between variables, like this is the case for our scenario parameters.
@@ 3 ESTIMATION OF GMCM PARAMETERS
x $\in\mathbb{V}^{d}=\bigl[F^{-1}_{1}(u_{1}),\,F^{-1}_{2}(u_{2}),\,\dots\,F^{-1}_{d}(u_{d})\bigr],$ (8b)
@@
$\mathcal{N}\!\Bigl(\sum_{k=1}^{K}\alpha^{k}\,\mu_{j}^{k}\;\big|\;0,\sigma\Bigr),$ (10)
$\mathcal{N}\!\Bigl(\sum_{k=1}^{K}\bigl[\alpha^{k}\bigl(\Sigma^{k}_{jj}+(\mu_{j}^{k})^{2}\bigr)\bigr]\big|\;1,\sigma\Bigr).$ (11)
@@ 4 EXPERIMENT
In this section, we present our experimental study of GMCMs for joint probability estimation of scenario parameters. We demonstrate that GMCMs surpass GCMs and perform better than, or at least comparably to, GMMs in accurately capturing the complex parameter dependencies. Our experiments conducted on real‐world datasets for certification relevant scenarios, confirm the previously reported advantages of GMCMs [19, 23]. They also extend these benefits to the domain of scenario parameter modeling.
@@ 4.1 REAL-WORLD SCENARIO DATASETS
    Deceleration: “the ‘other vehicle’ suddenly decelerates in front of the ‘ego vehicle”’[3, Annex 4, Appendix 3, para. 2.2(c)]
@@
For our investigation, we analyzed 13,613,530 and 4,111,855 real-world instances of the Follow Lead Vehicle Emergency Brake and Deceleration scenario, respectively. The data is derived from the Mercedes-Benz customer fleet, which collects information whenever its assistance systems detect a critical traffic situation, provided that customers have given their consent. In such cases, a full report describing the situation is sent to the backend for further analysis. For more details about the data used in this study, refer to [37].
@@ 4.2 RESULTS
In the second step, the MAP procedure for the estimation of the dependencies $\Theta$ among the scenario parameters was carried out using the Adam optimizer with the default learning rate of $10^{-3}$ [39]. The GMCMs were trained on both datasets with four mixture components. An automated selection using criteria such as the Bayesian Information Criterion (BIC) is proposed for future work (see Section 5).
To ensure a fair comparison, we matched the GMM component count to a comparable compute-time budget measured on the same machine. Specifically, we ran the four-component GMCM five times and used the median wall-clock time as the budget. We then chose the GMM component count so that its median wall-clock time (over five runs) matched this budget most closely. This procedure was applied independently to each dataset, resulting in GMMs with twelve components for the Follow Lead Vehicle Emergency Brake dataset and ten components for the Deceleration dataset. All experiments were run on a machine with an Intel Core i7-1365U (10 cores/12 threads) and 32 GB RAM running Windows 11. No GPU acceleration was used.
To illustrate the achieved expressive power of the GMCM, we visualize the bivariate marginal distribution of LeadVehicle Init HeadwayTime (s) and Ego InitSpeed Ve0 (kph) from the Follow Lead Vehicle Emergency Brake dataset via density heatmaps. Fig. 4 compares the original data distribution—which appears multimodal with non-Gaussian modes—with the estimated densities from the GMCM, the GMM, and the GCM. Visual inspection of the bivariate density heatmaps suggests that the GMCM most accurately captures the real data distribution.
@@
Unlike the heatmaps, which only allow model evaluation in two dimensions, this observation is quantitatively supported across the full-dimensional space by the Sinkhorn distance and the log-likelihood, where the GMCM demonstrates improved performance—most clearly in the Sinkhorn distance—while achieving a log-likelihood comparable to the GMM. In an analogous evaluation on the Deceleration scenario dataset using the same metrics, the GMCM performs comparably to the GMM, with the GMM holding a slight edge. The GCM performs markedly worse, as shown in Table 3.
Table 3: Log-likelihood, Sinkhorn distance, and median training wall-clock time achieved by GMCM, GCM, and GMM on our Follow Lead Vehicle Emergency Brake and Deceleration datasets. The best-performing model for each metric and dataset is shown in bold.
Model Log-Likelihood Sinkhorn Distance Wall-Clock Time
GMM -6.8610 0.542 1546 s
GCM -7.1706 0.511 41 s
GMCM -6.7923 0.436 1707 s
@@
Model Log-Likelihood Sinkhorn Distance Wall-Clock Time
GMM -9.3703 1.65 361 s
GCM -12.0305 11.8 15 s
GMCM -9.6185 1.69 289 s
@@
The Python implementation we used to achieve the results presented in this section is provided with the paper [42].
@@ 5 CONCLUSION
GMCMs demonstrated superior performance on one dataset and, on the other, performed comparably to GMMs (with a slight GMM edge) while exceeding the GCM, even without hyperparameter fine-tuning. This suggests that GMCMs offer a promising alternative to established approaches for probabilistic modeling of scenario parameters. Future work should explore the extent to which performance can be improved by tuning estimation parameters, such as selecting the optimal number of mixture components via the BIC and adjusting convergence criteria [44]. Moreover, GMCMs should be applied to other scenario types defined in UN R157, such as Cut-in and Cut-out. Their performance in higher-dimensional settings should also be evaluated, for example by incorporating parameters related to roadway and environmental conditions. Finally, we restricted our analysis to continuous parameters. How discrete scenario parameters could be integrated into the GMCM framework remains an open question and is left for future work.